### v2

# ArchCAD-400K: An Open Large-Scale Architectural CAD Dataset and New Baseline for Panoptic Symbol Spotting

Abstract
Recognizing symbols in architectural CAD drawings is critical for various advanced engineering applications. In this paper, we propose a novel CAD data annotation engine that leverages intrinsic attributes from systematically archived CAD drawings to automatically generate high-quality annotations, thus significantly reducing manual labeling efforts. Utilizing this engine, we construct ArchCAD-400K, a large-scale CAD dataset consisting of 413,062 chunks from 5538 highly standardized drawings, making it over 26 times larger than the largest existing CAD dataset. ArchCAD-400K boasts an extended drawing diversity and broader categories, offering line-grained annotations. Furthermore, we present a new baseline model for panoptic symbol spotting, termed Dual-Pathway Symbol Spotter (DPSS). It incorporates an adaptive fusion module to enhance primitive features with complementary image features, achieving state-of-the-art performance and enhanced robustness. Extensive experiments validate the effectiveness of DPSS, demonstrating the value of ArchCAD-400K and its potential to drive innovation in architectural design and construction.
1 Introduction
For a long time, CAD drawings have served as the universal language of architectural design, enabling seamless communication among designers, engineers, and construction personnel through standardized graphical primitives such as arcs, circles, and polylines. Accurate perception of these primitives in 2D CAD drawings is essential for various downstream applications, including automated drawing review and 3D building information modeling (BIM) [1, 2, 3, 4], as this ability enables the optimization of CAD-based workflows while enhancing efficiency and precision in architectural design and construction processes[5, 6].
Building upon earlier research in symbol spotting [7, 8, 9, 10], pioneering work formally defined the task of panoptic symbol spotting [11] for floor plan CAD drawings and established the benchmark dataset FloorPlanCAD [11, 12, 13, 14, 15, 16]. However, the manual annotation of line-grained labels is a highly time-consuming and labor-intensive process, severely limiting the dataset’s scale and diversity. This bottleneck restricts the ability to train models that can generalize across diverse building types, varying spatial scales, complex layouts, and a wide range of architectural component categories, thereby hindering their applicability to real-world scenarios.
Figure 1: Illustration of Layer-Block Structure. (a) Layers: A CAD drawing is organized as multiple overlaid layers, each serving a distinct functional or categorical purpose. (b) Blocks: Frequently used elements are defined as reusable blocks, allowing for efficient replication and modification. Best viewed in color.
(a) Raw CAD Drawings
(b) Panoptic labeled Drawings
Figure 2: An example of annotated drawings in ArchCAD-400K. Each geometric primitive is assigned both semantic and instance labels, with distinct semantic categories represented by different colors and instances visualized through convex hull masks. The chunks are extracted from complete drawings. Best viewed in color.
To address these limitations, we propose a highly efficient annotation pipeline for line-grained labels, which significantly reduces the annotation cost from 1,000 person-hours required for 16K data to 800 person-hours for 413K data. The core idea is to leverage the structured organization of floor plan drawings, including layers and blocks, to enable scalable and cost-effective large-scale annotation. Specifically, layers facilitate bulk annotation of specific categories through semantic aggregation (e.g., door and window layers), while blocks enable unified identification and annotation of repeated instances via an instance reuse mechanism. As shown in Figure 1, this layer-block structure, inherently designed to enforce drawing standards and ensure consistency in architectural documentation, forms the foundational framework for our automated annotation pipeline. In addition, for high data quality, we restrict data sources to completed drawings from leading design institutions and employ a fully vectorized annotation workflow to eliminate errors arising from image-to-vector conversions, with human experts reviewing and correcting automated results.
Based on the efficient and accurate data engine, we construct ArchCAD-400K, a large-scale floor plan CAD drawing dataset for the panoptic symbol spotting task. The dataset contains 5,538 complete drawings meticulously selected from 11,917 industry-standard drawings, with a total of 413,062 annotated chunks, surpassing FloorPlanCAD in scale by a factor of 26. While FloorPlanCAD primarily focuses on residential buildings, ArchCAD-400K covers a diverse range of building types, with residential structures accounting for only 14% and large-scale public and commercial facilities forming the majority. The average area of drawings in ArchCAD-400K spans 11,000 m2, significantly larger than FloorPlanCAD’s 1,000 m2 average, with 51.9% ranging from 1,000 to 10,000 m2 and 4.4% exceeding 100,000 m2. Furthermore, ArchCAD-400K introduces a comprehensive semantic categorization, including 27 categories such as structural components (e.g., columns, beams), non-structural elements (e.g., doors, windows), and drawing notations (e.g., axis lines, labels), with 14 categories each containing over 1 million primitives. This extensive scale, diversity, and detailed annotation make ArchCAD-400K a robust resource for advancing AI models in construction industry. An example from ArchCAD-400K is illustrated in Figure 2.
We further propose a novel framework for panoptic symbol spotting, named Dual-Pathway Symbol Spotter (DPSS), which incorporates an adaptive fusion module to effectively enhance primitive features with complementary image features. This design enables the model to achieve superior performance and enhanced robustness compared to existing methods while demonstrating strong scalability and generalization capabilities on larger-scale datasets. In summary, our work presents the following contributions:
(1) We develop a highly efficient annotation pipeline specifically designed for floor plan CAD drawings, which generates high-quality annotations with improved efficiency compared to image-based manual annotation, significantly reducing the annotation cost from 1,000 person-hours for 16K data to 800 person-hours for 413K data.
(2) We introduce ArchCAD-400K, a large-scale floor plan CAD drawing dataset that surpasses the current largest FloorPlanCAD dataset by an order of magnitude in size and exhibits greater diversity in terms of building types, spatial scales, and architectural component categories.
(3) We propose DPSS, a novel framework for panoptic symbol spotting that achieves state-of-the-art performance on FloorPlanCAD and ArchCAD-400K, surpassing the second-best method by 3% and 10%, respectively, with exceptional accuracy, robustness, and scalability.
2 Related Works
2.1 Floor Plan Datasets
Several datasets have been developed for floor plan analysis. SESYD [9] comprises 1,000 synthetic vectorized documents with ground truth annotations. FPLAN-POLY [10] contains 42 floor plans derived from images [17] for spatial relationship analysis. Cubicasa [3] provides 5,000 floor plan images annotated with over 80 object categories, focusing on residential layouts. RFP [18] includes 2,000 annotated floor plans with detailed room-level information for Asian residential buildings. FloorPlanCAD [11] offers 16,103 vector-graphic floor plans in “.svg" format, annotated across 35 object categories. LS-CAD [16] introduces a test set of 50 full-size CAD drawings with an average area of 1,000 square meters. However, these datasets are limited in scale and rely on labor-intensive annotation, hindering large-scale training.
2.2 Large-scale Vision Datasets
Large-scale vision datasets have significantly advanced deep learning models for complex visual understanding tasks[19, 20, 21, 22, 23, 24, 25, 26, 27], including floor plan recognition. Image-based datasets such as ImageNet [19] (14M images, 21K categories) and COCO [20] (300K images with object detection, segmentation, and captioning labels) have been instrumental in developing state-of-the-art models for object recognition, classification, and scene understanding. These datasets provide rich annotations that enhance model generalization, improving performance across general and domain-specific tasks. Nevertheless, large-scale vector-based datasets and efficient annotation processes for floor plan understanding remain scarce.
2.3 Panoptic Symbol Spotting
The panoptic symbol spotting task, initially proposed in [11], involves the simultaneous detection and classification of architectural symbols (e.g., doors, windows, stairs) in floor plan CAD drawings. While traditional methods [7] focus on countable instances (e.g., windows, tabls), Fan et al. [11] extended this to uncountable objects (e.g., walls, railings), inspired by [28]. They introduced PanCADNet, which integrates Faster R-CNN [29] for countable instances and Graph Convolutional Networks [30] for uncountable elements. Subsequently, Fan et al. [12] proposed CADTransformer, utilizing HRNetV2-W48 [31] and Vision Transformers [32] for primitive tokenization and embedding aggregation. Zheng et al. [15] adopted graph-based representations with Graph Attention Networks for semantic and instance-level predictions. Liu et al. [14] introduced SymPoint, exploring point set representations, later enhanced by SymPointV2 [13] through layer feature encoding and position-guided training. In contrast, CADSpotting [16] densely samples points along primitives and employs Sliding Window Aggregation for efficient panoptic segmentation of large-scale CAD drawings. While these methods perform well on FloorPlanCAD, they struggle on our more diverse, complex, and large-scale dataset, highlighting the need for robust and scalable solutions.
Figure 3: Overall pipeline of the annotation process. Primitives are labeled by automated methods followed by manual modification. Best viewed in color.
Figure 4: Annotation format of the ArchCAD-400K
3 Annotation pipeline of ArchCAD-400K
We carefully gathered over 11917 complete CAD drawings from the industry, covering a wide range of building types. From the initial selected drawings, 5,538 drawings were validated as strictly conforming to the layer-block organizational standards. Through automated annotation combined with expert refinement, we generated approximately 413,062 chunks with line-grained annotations, each sized at 14m×14m, aligning with the design of FloorPlanCAD [11]. The overall pipeline of the annotation process is shown in Figure 3. In the following, we outline the annotation process of ArchCAD-400K in depth.
3.1 Annotation Format
The ArchCAD-400K employs a structured annotation format tailored for the panoptic symbol spotting task, as defined in [11]. Each graphical primitive (e.g., line, arc, circle) in the drawing is characterized by a dual-identifier pair, $(l_{k},z_{k})$, where $l_{k}$ denotes its semantic category and $z_{k}$ represents the instance identifier. Primitives sharing the same $z_{k}$ value are considered to be part of the same instance. Figure 4 illustrates the annotation format for architectural layouts. This format uses an SVG file structure, representing elements such as paths and circles. Each graphical primitive is associated with a semantic label, an instance identifier, and an RGB color for visualization.
3.2 Layer-Block Standardization Screening
Manual annotation of floor plans is inefficient due to crowded layouts and inconsistent symbols. Standardized drawings use layers for bulk annotation (e.g., door layer) and blocks for identifying repeated instances, clarifying topological and semantic relationships for automation.
However, the feasibility of automated annotation depends on standardized drawings. Non-standard drawings can have ambiguous layer names and mixed primitives, leading to semantic confusion. Notably, professional design institutes typically maintain internal layering standards, which can be normalized and processed to establish machine-interpretable structures. Thus, we restrict our data to completed drawings from leading design institutions and apply a layer naming validation algorithm against a reference table, discarding drawings with over 5% deviation. This dual quality control ensures standardization of the input data, providing a solid foundation for future automated annotation based on layer-block standardization. The reference table can be found in our supplementary materials.
3.3 Automated Annotation with Expert Refinement
The automated annotation process, based on the layer-block standardization, significantly reduces the workload of manual annotation, providing an efficient foundation for large-scale drawing annotation. However, despite the substantial efficiency gains, the accuracy of automated annotation remains limited, particularly when handling non-standard drawings, which inevitably persist despite the standardization screening. Semantic ambiguities and mixed graphic primitives within layers may lead to annotation errors, making human intervention essential to ensure accuracy.
To address this, we engaged 10 professional architectural drafters with over 3 years of experience to refine the automatically generated annotations using an intuitive vector-based interface. Compared to the raster-based annotation method [11], such as polygon or bounding box labeling, directly operating on vector primitives not only eliminates potential errors introduced by raster-to-vector conversions but also overcomes the challenge of distinguishing overlapping instances in raster images, thereby reducing potential ambiguities. We created tools that automatically check CAD drawings against layer-block rules. Errors are highlighted directly on the drawings, allowing human corrections within seconds per file.
To achieve line-grained panoramic annotation, we have adopted a two-stage annotation system, as shown in Figure 3. In the semantic annotation stage, the system first classifies the layers based on their names and content into predefined semantic categories and displays the categories of each graphic primitive in professional CAD drawing software, allowing experts to directly correct inaccuracies through a vector-based operation interface. In the instance annotation stage, specific semantic categories are extracted and processed individually, isolating instances of different categories to reduce interference from surrounding information. The system automatically generates instance mask proposals based on existing block information in the drawings, which are then reviewed by experts who manually correct any errors. Through this two-stage annotation system, we completed the panoramic annotation of 5,538 drawings in approximately 800 hours, achieving an efficiency more than 10 times that of the FloorPlanCAD dataset. This process not only significantly enhances annotation efficiency but also ensures high accuracy and reliability.
4 Exploring ArchCAD-400K
Dataset Source Scale Data Format Annotation Elements
Size Total Area Raster Vector Type Method Architectural Structural Notation
FPLAN-POLY [10] Internet 48 <$5\times 10^{4}$ m2 Instance Human
SESYD [9] Synthetic 1000 <$5\times 10^{5}$ m2 Instance Human
FloorPlanCAD [11] Industry 16K $1.6\times 10^{6}$ m2 Panoptic Human
ArchCAD-400K Industry 413K $8.1\times 10^{7}$ m2 Panoptic Auto & Human
Table 1: Comparison of ArchCAD-400K to existing CAD drawing datasets.
We compare our ArchCAD-400K with existing datasets, as summarized in Table 1. Publicly available datasets, such as SESYD [9], FPLAN-POLY [10], and FloorPlanCAD [11], exhibit limitations in terms of source, scale, data format, annotation type and method, as well as the range of elements included in floor plans. These constraints affect their applicability to real-world architectural analysis.
4.1 Extended Building Diversity
Figure 5: (a) Different project types of drawings; (b) Area distribution of drawings.
Figure 6: Number of annotated primitives for 27 semantic classes. Bar colors match each class’s visualization.
Previous floor plan datasets predominantly originated from residential buildings, with a primary emphasis on indoor layouts. This limitation constrained their applicability in analyzing a wider variety of project types. In contrast, our dataset is curated from a more diverse array of buildings, as illustrated in Figure 5, with residential structures comprising only 14% of the total. A substantial portion consists of large-scale public and commercial buildings, including office complexes, industrial parks, and other expansive facilities. This diverse composition provides a richer and more representative architectural dataset, enabling applications across a wider variety of building scenarios.
4.2 Larger Data and Spatial Scale
ArchCAD-400K stands out for its unprecedented scale, comprising 5,538 complete drawings and 413,062 chunks, surpassing FloorPlanCAD’s 15,663 chunks by over 26 times. This substantial scale significantly enhances the robustness and generalization capabilities of AI models for floor plan analysis.
In addition to the data scale, ArchCAD-400K features much larger drawings compared to existing datasets, which are generally restricted to small areas under 1,000 m2. The average drawing in our dataset spans 11,000 m2, closely reflecting real-world architectural dimensions. As shown in Figure 5, 51.9% of the drawings cover areas between 1,000 and 10,000,m2, while 4.4% exceed 100,000 m2, representing a wide spectrum of building areas and configurations.
4.3 Diverse Element Types
Traditional floor plan datasets typically focus on common non-structural elements, such as furniture, equipment, and decorative features. While such categorizations are sufficient for residential interior layouts, they often fail to capture critical components when applied to more diverse building types, particularly in complex architectural or industrial settings. To address this limitation, our ArchCAD-400K adopts a comprehensive semantic categorization that extends beyond traditional architectural elements. Specifically, we introduce annotations for structural components (e.g., columns, beams, and holes) and drawing notations (e.g., axis lines, labels, and markers) in addition to non-structural elements. Structural components are fundamental to reconstructing the overall building structure as they define the load-bearing framework and spatial organization of the architectural design. In contrast, drawing notations are ubiquitous in real-world drawings and play a critical role in ensuring accurate interpretation and practical implementation. This tripartite categorization, encompassing non-structural, structural, and notation elements, enables a broader range of applications, including indoor navigation, architectural design, and structural analysis.
ArchCAD-400K provides detailed annotations for 27 categories, with the distribution of primitives across these categories visually represented in Figure 6. Notably, 14 of these categories each encompass more than 1 million primitives, underscoring the extensive scale of our dataset. Additionally, our dataset includes 7 countable instance categories. While this count is fewer than that of FloorPlanCAD, our dataset surpasses it in terms of diversity and complexity within each category, presenting a more formidable challenge. For instance, we consolidate various fine-grained furniture types from FloorPlanCAD, such as chair, table, and bed, into a single unified “furniture" category. Visual examples of the categories can be found in our supplementary materials.
5 Method
(a) DPSS Framework
(b) Adaptive fusion
Figure 7: The overall architecture of the proposed Dual-Pathway Symbol Spotter (DPSS). CAD drawings are encoded through both point backbone and image backbone. Features from dual pathways are aligned and fused in adaptive fusion module, then pass through a transformer decoder to generate mask predictions and classifications at primitive level. Best viewed in color.
To advance panoptic symbol spotting on large-scale datasets, we propose Dual-Pathway Symbol Spotter (DPSS), a novel framework designed for enhanced robustness and generalization. Unlike previous methods that rely on prior knowledge, such as color and layer information, DPSS eliminates these dependencies, improving adaptability across diverse datasets. As illustrated in Figure 7, the approach incorporates a dual-path feature extraction module, consisting of an image branch and a point cloud branch. The image branch encodes rendered vector graphics to capture global contextual features, while the point cloud branch represents primitives as points to extract fine-grained geometric details. An adaptive fusion mechanism aligns and integrates these multimodal features, ensuring effective feature representation for each primitive. Finally, a transformer decoder performs panoptic segmentation of the entire graphic, enabling precise symbol spotting in complex drawings.
5.1 Two-stream Primitive Encoding Module
The extraction of high-quality features for vector graphic panoramic segmentation is crucial as it directly affects segmentation performance. Unlike raster images, vector graphics consist of primitives, yet research on their modeling remains limited, highlighting the need for better feature extraction methods. CADTransformer [12] rasterizes vector graphics and samples features at the primitive centers, while SymPoint [14] encodes primitives as points using geometric attributes and a point cloud encoder[33]. Image-based approaches capture semantics well but struggle with fine-grained tasks, whereas point-based methods handle instances effectively but are less suited for sparse structures like walls. To address this, we combine visual and geometric encoding. Our visual encoder, based on HRNetV2 [34], extracts feature maps from rendered images. Meanwhile, a Point Transformer-inspired point cloud encoder generates geometric features for each primitive, ensuring a more comprehensive representation.
5.2 Adaptive Fusion Module
We use an image sampler to extract semantic features for each primitive. Given the primitive center $(x_{c},y_{c})$ and the feature map $F\in\mathbb{R}^{H\times W\times C}$, the sampler produces point-wise image features $V$. Since sampling points may fall between pixels, bilinear interpolation is applied:
$V_{p}=\mathcal{K}(F_{N(p^{\prime})}),$ (1)
where $V_{p}$ is the extracted feature, $\mathcal{K}$ is the interpolation function, and $F_{N(p^{\prime})}$ denotes neighboring pixels. The combination of primitive and image features is challenging due to size variations, noise, and overlapping. To address this, we introduce a geometry-guided fusion layer. First, graphic features $X_{p}$ and image features $V_{p}$ are transformed via an MLP, then concatenated. A weight matrix $w$ is computed as:
$w=\sigma(W_{1}{\rm tanh}(W_{2}X_{p}+W_{3}V_{p})),$ (2)
where $W_{1}$, $W_{2}$, $W_{3}$ are learnable parameters. Finally, the weighted features yield the final representation:
$U_{p}={\rm concat}(X_{p},wV_{p}).$ (3)
5.3 Decoder and Loss Function
The decoder is designed based on the architectures of DETR [35] and Mask2Former [36], achieving panoptic symbol spotting by predicting primitive-level masks along with their categories. The loss function is composed of a standard cross-entropy loss ($L_{cls}$) for class predictions, a binary cross-entropy loss ($L_{bce}$), and a Dice loss ($L_{dice}$[37] for mask predictions. The overall loss is formulated as a weighted sum of the three losses $L=\lambda_{cls}L_{cls}+\lambda_{bce}L_{bce}+\lambda_{dice}L_{dice}$, where $\lambda_{cls}$, $\lambda_{bce}$, and $\lambda_{dice}$ denote the weight for each loss term respectively.
6 Experiments
We split our ArchCAD-400K dataset into training, validation, and test sets using a 7:1:2 ratio, ensuring that each drawing and its corresponding annotations appear in only one split. This results in 289,144 annotated samples for training, 41,306 for validation, and 82,612 for testing. Following the definition of panoptic symbol spotting, we evaluate the model performance using Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ). The formulation of these metrics can be found in [11]. To evaluate model performance on ArchCAD-400K, we compare existing methods and our proposed method DPSS, with detailed results presented below.
6.1 Implement Details
We adopt HRNetW48 [31] pretrained on COCO-Stuff [38] as the backbone for the visual branch. The point cloud branch uses the PointTransformerV2 [39] encoder pretrained on ScanNetV2 [40]. The model is trained on 8 NVIDIA A800 GPUs and the optimizer is AdamW. On FloorplanCAD [11] dataset, we set a batch size of 2 per GPU and with a learning rate of $2\times 10^{-4}$ and a weight decay of 0.1. The model is trained for 50 epochs. On ArchCAD-400K, we set a batch size of 4 per GPU and train the model for 10 epochs with a learning rate of $2\times 10^{-4}$ and the same weight decay. Other baseline methods on ArchCAD-400K are trained for 10 epochs under their default configurations.
6.2 Quantitative Evaluation
Method Additional prior inputs Total Thing Stuff
PQ SQ RQ PQ SQ RQ PQ SQ RQ
SymPointV2[13] w/ 90.1 96.3 93.6 90.8 96.6 94.0 80.8 90.9 88.9
CADSpotting[16] w/ 88.9 95.6 93.0 89.7 96.2 93.2 80.6 89.7 89.8
DPSS w/ 89.5 96.2 93.1 90.4 96.6 93.5 79.7 91.1 87.5
CADTransformer[12] w/o 68.9 88.3 73.3 78.5 94.0 83.5 58.6 81.9 71.5
GAT-CADNet[15] w/o 73.7 91.4 80.7
SymPoint[14] w/o 83.3 91.4 91.1 84.1 94.7 88.8 48.2 69.5 69.4
SymPointV2[13] w/o 83.2 91.3 91.1 85.8 92.5 92.7 49.3 70.3 70.1
DPSS w/o 86.2 93.0 92.6 88.0 94.1 93.5 64.7 83.0 77.9
Table 2: Panoptic symbol spotting results on FloorplanCAD [11] dataset
Method Additional prior inputs Semantic spotting Instance spotting
F1 wF1 AP50 AP75 mAP
SymPointV2 [13] w/ 89.5 88.3 71.3 60.7 60.1
CADSpotting [16] w/ 93.5 93.9 72.2 69.1 69.0
DPSS w/ 93.1 93.2 74.8 71.0 70.8
DeepLabv3+R101 [41] w/o 68.8 71.4
DINO [42] w/o 64.9 54.9 47.5
CADTransformer [11] w/o 82.2 80.1
SymPoint [14] w/o 86.8 85.5 66.3 55.7 52.8
SymPointV2 [13] w/o 87.0 86.3 66.4 57.7 57.5
DPSS w/o 92.0 93.2 67.0 61.8 61.5
Table 3: Semantic and instance spotting results on FloorPlanCAD
Quantitative comparison on FloorPlanCAD. We evaluate the performance of various methods on the FloorPlanCAD dataset, as shown in Table 2 and Table 3, focusing on both scenarios: one with prior information (such as layers and color) and one without. This dual evaluation is motivated by the intrinsic properties of CAD drawings, where layer and color information naturally encode semantic relationships among graphical primitives. For instance, primitives within the same layer or sharing similar colors are more likely to belong to the same semantic category.
Notably, SymPointV2 extends SymPointV1 by incorporating layer feature encoding, while CADSpotting relies on dense sampling of primitives based on color and positional information. The experimental results in Table 2 show that our DPSS performs similarly to these methods when layer and color information is available. However, in the absence of such priors, SymPointV2 experiences a significant performance drop, while DPSS remains stable, achieving a 3% higher total PQ than the second-best method. Particularly, DPSS outperforms SymPointV2 by 15% in PQ for the "Stuff" category, highlighting its robustness and generalization capabilities. As shown in Table 3, when layer and color information is unavailable, DPSS significantly surpasses the second-best method in both Semantic Spotting and Instance Spotting, further underscoring its superior performance across various metrics.
While our results demonstrate strong performance with the inclusion of prior information, we place greater emphasis on performance without these additional priors. This is because the absence of layer and color data more closely reflects real-world conditions, where such information is not often available.
Quantitative comparison on ArchCAD-400K.
Method Total Thing Stuff
PQ SQ RQ PQ SQ RQ PQ SQ RQ
CADTransformer[12] 60.0 89.7 66.9 52.5 83.6 62.7 70.1 96.7 72.5
SymPoint[14] 47.6 86.1 55.3 51.4 91.9 55.9 39.9 73.9 54.0
SymPointV2[13] 60.5 88.0 68.8 62.4 91.7 68.1 52.8 73.7 71.7
DPSS 70.6 90.2 78.2 65.6 92.4 70.9 77.6 87.8 88.4
Table 4: Panoptic symbol spotting results on ArchCAD-400K
Method Semantic spotting Instance spotting
F1 wF1 AP50 AP75 mAP
CADTransformer[12] 84.1 83.4
SymPoint[14] 76.8 62.0 36.1 30.9 30.8
SymPointV2[13] 69.8 69.3 44.9 40.2 39.7
DPSS 87.8 84.1 45.6 41.1 40.7
Table 5: Semantic and instance spotting results on ArchCAD-400K
We also evaluate the performance of various methods on our ArchCAD-400K dataset, as shown in Table 4 and Table 5, comparing DPSS with other methods, including CADTransformer, SymPoint, and SymPointV2. Importantly, the results presented here are based on the scenario where layer and color prior information is not utilized, reflecting more realistic conditions.
ArchCAD-400K, being more challenging than FloorPlanCAD, presents additional complexities, which is reflected in the generally lower performance across all methods. For example, SymPointV2, which achieves a total PQ of 83.2% on FloorPlanCAD, shows a significant drop in performance on ArchCAD-400K, with the total PQ dropping by more than 22%, down to 60.5%.
For semantic and instance spotting, DPSS shows significant improvements over other methods, particularly in terms of scalability and robustness. In semantic spotting, DPSS achieves an F1 score of 87.8%, outperforming the CADTransformer (84.1%) and SymPoint (76.8%), and showing a substantial gain of nearly 18% over SymPointV2 (69.8%). In instance spotting, DPSS achieves an mAP of 40.7%, outpacing SymPointV2 (39.7%) and SymPoint (30.8%). For panoptic symbol spotting, DPSS also leads the field, achieving a total PQ of 70.6%, notably higher than SymPointV2 (60.5%) and SymPoint (47.6%). DPSS excels particularly in the Stuff category, where it achieves a Stuff PQ of 77.6%, representing a 24.8% improvement over SymPointV2. These results underscore DPSS’s robustness and ability to generalize effectively, especially in challenging categories that are harder to define.
In summary, even without additional prior information, DPSS demonstrates superior performance across all major metrics compared to other methods, reflecting its strong scalability and robustness. Our ArchCAD-400K dataset, being more challenging, shows that DPSS is not only capable of achieving high performance in more complex scenarios, but also adaptable to real-world applications where such prior information is seldom accessible. These findings position DPSS as a leading method for panoptic symbol spotting in architectural CAD drawings.
6.3 Dataset Ablations
Generalization ability of ArchCAD-400K. To validate the generalization ability of ArchCAD-400K, We implement DPSS on FloorPlanCAD[11] dataset with a dual-pathway encoder pretrained on ArchCAD-400K. Results are illustrated in Figure 8. The pre-trained model achieves faster and more stable convergence, surpassing the performance of non-pre-trained models at 50 epochs within 30 epochs. It suggests that ArchCAD-400K covers FloorPlanCAD and furthermore has more diverse patterns and properties for models to generalize across datasets.
Scaling performance of ArchCAD-400K. We conducted experiments to validate the scaling law in the context of panoptic symbol spotting. Using a subset of our ArchCAD-400K and adopting DPSS as the baseline network, we evaluated the training performance across different data scales, ranging from 10K samples to the full 400K samples. As demonstrated in Figure 8 and 8, within 10K to 400K data range, doubling the dataset size consistently reduces the loss by 2.33 and improves the PQ metric by 6.40% in average. The results indicate that, for panoptic symbol spotting tasks, large-scale datasets significantly enhance model performance.
Figure 8: (a) Comparison of model performance: pretrained vs. non-pre-trained on ArchCAD-400K; (b) Loss convergence across different dataset sizes; (c) Panoptic quality across different dataset sizes.
6.4 Qualitative Results
The qualitative results on our ArchCAD-400K are illustrated in Figure 9. In the presented cases, DPSS demonstrates superior performance than SymPointV2 [13]. In architectural diagrams (row 1), characterized by a diversity of component types and numerous interfering lines, DPSS is capable of filtering out such disturbances to achieve precise spotting results. In structural drawings (row 2), where the shapes of components exhibit high similarity, DPSS effectively integrates contextual information to discern semantic differences between analogous components. More qualitative results on ArchCAD-400K can be found in our supplementary materials.
(a) GT
(b) DPSS
(c) SymPointV2 [13]
Figure 9: Qualitative comparison between DPSS and SymPointV2 on ArchCAD-400K. (a): Ground truth. (b): Results from DPSS. (c): Results from SymPointV2. Best viewed in color.
7 Conclusion
In this work, we address the challenges in panoptic symbol spotting for architectural CAD drawings by introducing an efficient annotation pipeline, a large-scale dataset (ArchCAD-400K), and a novel framework (DPSS). Our pipeline reduces annotation costs, enabling the creation of ArchCAD-400K, which surpasses existing datasets in scale and diversity. With 413,062 annotated chunks from 5,538 drawings, ArchCAD-400K covers diverse building types and spatial scales, advancing AI models in architectural design. DPSS, equipped with an adaptive fusion module to effectively enhance primitive features with complementary image features, achieves state-of-the-art performance on FloorPlanCAD and ArchCAD-400K, demonstrating superior accuracy, robustness, and scalability. We hope that our ArchCAD-400K will catalyze further progress in this domain.
Appendix A Detailed description of categorization
In Figure 10, we present some visual examples of content from each category. Engineering symbols from different categories can share great similarity. For example, simple rectangles can represent columns, holes, foundation, or furniture. Similarly, pairs of parallel lines might denote pipeline, beam, or wall. At the same time, some instance can be drawn in diverse forms. As illustrated in Figure 10, there are at least six different ways to represent a door, and the appearance of stairs can vary much in different scenarios.
Figure 10: Visual examples of content in each category.
Appendix B Additional information about the annotation pipeline
Our annotation pipeline adopts regex matching to map layer names to semantic classes. The layer names in standardized drawings have a hierarchical format like [Discipline]-[Category]-[Modifier]. In Table 6, we list part of the standardized naming table for doors, walls, and stairs. Categories can be matched through the keyword in the table.
Table 6: Examples of correspondence between layer names and semantics
Semantic label Standard layer specifications Content description
Door A-DOOR The surface structure of the door.
A-DOOR-FRAM The internal steel frame supporting the door.
A-DOOR-HEAD The line indicating the position of the door beam.
A-DOOR-ROLL The layout or design of the fireproof roller shutter.
Wall A-WALL-BLOK Masonry block wall for structural stability.
A-WALL-CONC Concrete wall with high strength and fire resistance.
A-WALL-STUD Lightweight partition with stud framing and drywall.
A-WALL-PRHT Partial-height wall for spatial division.
A-WALL-SCRN Metal wall for electromagnetic shielding.
A-WALL-FINI Final surface treatment or cladding of a wall.
A-WALL-INSU Insulation layer for thermal or acoustic performance.
A-WALL-TPTN Partial-height wall for restroom privacy.
A-WALL-EXPL Reinforced wall to withstand blast pressures.
S-WALL-LINE The outline or framework of a wall in structural drawings.
S-WALL-HATCH Represents the filling or material pattern of a wall in structural drawings.
Stairs A-STRS-TREA Stepping surface and vertical connector in stair construction.
A-STRS-ESCL Mechanical moving stairs for vertical transportation between floors.
A-STRS-HRAL Safety rails along stair edges for fall prevention.
S-STRS-LINE Outline representing stair geometry in structural drawings.
Appendix C Additional Examples of ArchCAD-400K
Figure 2 shows an example of well-labeled large architectural drawings. We further show two labeled drawings in our dataset in Figure 11. Each of them covers an area over 8000 m2, with various types of components.
(a) Raw CAD Drawings
(b) Panoptic labeled Drawings
Figure 11: Additional example of the annotated drawings.
Appendix D Additional Quantitative Evaluations
We present the detailed experimental results of various methods on the ArchCAD-400K, including the panoptic quality (PQ) for each category as well as the mean IoU (Intersection over Union) for each category. The IoU for each category is obtained by calculating the intersection over union between the predicted panoptic segmentation masks and the ground truth masks. Different methods excel at handling different types of objects, DPSS achieves higher PQ and IoU metrics. For the current spotting results, there is still significant room for improvement.
Table 7: Quantitative results for panoptic symbol spotting of each category on ArchCAD-400K
class DPSS CADTransformer [11] SymPoint [14] SymPointV2 [13]
IoU PQ IoU PQ IoU PQ IoU PQ
symbol 84.4 89.8 62.4 70.7 76.4 80.9 74.2 79.3
axis 90.8 85.4 23.2 32.9 58.3 56.2 71.8 83.7
door 64.4 75.9 53.2 64.9 63.9 75.6 82.6 79.1
floor slab 40.8 85.6 14.4 80.3 20.9 82.8 47.7 69.4
elevator 45.6 77.6 26.5 63.9 54.6 73.3 28.0 69.2
stairs 60.4 65.4 35.3 45.2 41.3 55.4 43.7 78.0
furniture 72.5 88.7 66.5 73.3 63.2 83.9 43.4 54.0
hole 61.1 70.2 32.1 40.0 59.6 53.8 50.2 60.7
window 47.8 57.2 52.4 55.0 64.7 68.4 57.8 70.8
curtain wall 65.9 83.7 48.8 61.4 50.3 72.6 33.9 39.1
wall 73.4 75.6 39.1 45.9 52.7 57.7 42.3 70.7
concrete column 65.8 83.6 59.5 66.8 64.4 71.6 63.9 69.8
steel column 48.6 80.3 50.7 71.8 52.9 82.8 51.1 79.2
concrete beam 81.8 81.7 43.8 45.0 61.5 62.1 48.3 77.9
steel beam 79.1 73.0 19.7 34.9 47.5 52.2 71.5 77.4
parking spot 66.8 73.2 57.9 76.9 46.0 79.3 72.1 69.0
roof 3.5 28.7 3.7 39.1 5.7 50.1 55.6 76.1
base 85.0 90.3 65.8 75.2 71.2 83.1 16.3 21.3
bracing 19.6 33.8 7.8 20.0 20.2 23.2 67.6 81.1
rebar 72.7 93.1 39.4 69.4 56.5 81.1 53.5 75.9
equipment 37.2 45.8 1.0 7.0 30.8 31.8 28.4 29.1
handrail 73.6 65.9 31.2 31.0 40.2 42.7 70.3 85.8
pipe 44.4 48.3 32.9 34.4 51.7 41.7 52.5 42.4
window curtain 50.4 80.3 58.6 72.1 49.2 67.7 46.3 47.5
construction joint 0.0 5.3 0.0 0.3 2.5 4.8 4.1 3.7
embedment 0.0 0.0 0.0 0.0 0.0 0.0 1.2 2.8
hydrant 70.7 71.7 68.6 64.5 74.1 69.4 49.5 34.0
overall 70.6 67.04 47.6 49.1 60.5 59.1 60.0 60.8
Appendix E Additional Qualitative Evaluations
Qualitative results on FloorPlanCAD [11] are shown in Figure 12, DPSS performs better despite the noise of the filler graphics and overlaps between instances.
(a) GT
(b) DPSS
(c) SymPointV2[13]
(d) GT
(e) DPSS
(f) SymPointV2[13]
Figure 12: Additional qualitative comparison between DPSS and SymPointV2 on ArchCAD-400K. (a) and (d): Panoptic annotations. (b) and (e): Results from DPSS. (c) and (f): Results from SymPointV2.